Trimmed the lower branches from the north poplar tree to allow the motion sensor light to trigger on motion in the north parking lot.
Installed ‘Emergency Exit’ signage on the north classroom door.
Installed and tested an exterior motion sensor light on the north classroom wall.
Terra 2 classroom computer:
– Updated Real Player, DIVX+, and ITunes software.
– Replaced keyboard.
– Adjusted sound quality.
– Applied weed killer to the viewing deck.
– Eliminated a wasp nest.
– Installed IP9258 power bar and monitor at Tatla Lake.
– Performed annual maintenance on AllSky cameras and computers at Mile 108 and Tatla Lake.
– North ramp has been backfilled and compacted. Awaiting railings.
– Placed 6 spare patio blocks on the north ramp landing.
– Dome has been pressure washed, taped, sealed, and painted.
– Dome shutters have been painted.
– New flooring was installed in the dome room, stairwell, dome hallway, and control room.
– Replaced the furnace filter.
– Reattached the edging laminate to several classroom tables.
– Weatherstripped the north classroom door.
– Cleaned and restocked the bathroom.
– Repaired one section of the west soffit.
– Scraped and painted the south gable end and soffit.
– Reinstalled missing flashing on the large dome shutter.
– Removed excess material from the classroom to the basement.
– Swept and washed the classroom and hallway floors.
Many, many volunteers!
– Filled the driveway and parking lot ruts and potholes with donated road crush.
– Pulled weeds from the viewing deck.
– Did a general cleanup of the observatory grounds.
– The new classroom computer is now functional.
– Replaced the motherboard capacitors in two of the meteor monitoring computers.
– Wireless handpad has been installed and tested.
– Autotracking wiring from the ST10X to handpad control box has been installed and terminated.